# HebID: Detecting Social Identities in Hebrew-language Political Text

### v1

@@ Abstract
Political language is deeply intertwined with social identities. While social identities are often shaped by specific cultural contexts and expressed through particular uses of language, existing datasets for group and identity detection are predominantly English-centric, single-label and focus on coarse identity categories. We introduce HebID, the first multilabel Hebrew corpus for social identity detection: 5,536 sentences from Israeli politicians’ Facebook posts (Dec 2018–Apr 2021), manually annotated for twelve nuanced social identities (e.g., Rightist, Ultra-Orthodox, Socially-oriented) grounded by survey data. We benchmark multilabel and single-label encoders alongside 2B–9B-parameter seq2seq LLMs, finding that Hebrew-tuned LLMs provide the best results (macro- = 0.74). We apply our classifier to politicians’ Facebook posts and parliamentary speeches, evaluating differences in popularity, temporal trends, clustering patterns, and gender-related variations in identity expression. We utilize identity choices from a national public survey, enabling a comparison between identities portrayed in elite discourse and the public’s identity priorities. HebID provides a comprehensive foundation for studying social identities in Hebrew and can serve as a model for similar research in other non-English political contexts.111https://github.com/guymorlan/hebid/
@@ 1 Introduction
Social identities—such as political ideology, religious affiliation, or demographic group membership—are powerful drivers of political behavior and public discourse (Tajfel and Turner, 1979a, b). Yet existing NLP resources for identity detection remain almost entirely English-focused, single-label, and rely on coarse categories (e.g., party or ethnicity). In this work, we introduce HebID, the first publicly released Hebrew dataset for fine-grained, multilabel social identity detection in political text, grounded in both domain expertise and large-scale survey evidence. We utilized 12 waves of a national survey (N = 1,769) each of which included questions designed to identify the twelve expert-defined social identities most salient to Israeli citizens (e.g., Rightist, Ultra-Orthodox, Socially-oriented). These survey-derived categories were then used to guide the annotation of 5,536 sentences sampled from Facebook posts by Israeli politicians (December 2018–April 2021), ensuring that our labels reflect real-world identity salience.
We benchmark a suite of modern architectures on the Facebook data—(i) multilabel encoder models; (ii) single-label encoder models; and (iii) 2B–9B-parameter seq2seq LLMs, finding that Hebrew-tuned seq2seq models (specifically, DictaLM2.0) achieve the highest macro-F1 (0.743), outperforming encoder-only baselines by over six points. We then assess cross-genre generalization by applying our best model to 500 Knesset speech excerpts, achieving a comparable macro-F1 of 0.72.
Finally, we link three complementary sources in our analysis of how identities behave: politicians’ Facebook posts, parliamentary speeches from the Knesset floor, and survey responses that showcase public identification. We (1) compare identity prevalence and correlations between social media, parliamentary speech, and the public; (2) document how identity discourse surges around election cycles; (3) uncover coherent “bundles” of co-occurring identities; and (4) quantify gender-related variation in identity expression.
@@
Dataset: a multilabel Hebrew corpus of 5,536 sentences annotated for 12 social identities, empirically grounded in survey data.
Benchmarks: comprehensive comparison of encoder-only and seq2seq LLMs, showing the superiority of the latter.
Cross-genre evaluation: demonstration of model generalization to parliamentary speech.
Sociolinguistic analysis: novel insights into identity popularity, temporal dynamics, bundling, and gendered expression across multiple sources.
@@ 2 Related Work
Early work extended named-entity recognition to capture social groups as entities. Zanotto et al. (2024) introduce GRIT, annotating Italian news and parliamentary text for spans referring to demographic, national, or partisan groups, and fine-tune BERT to identify them. Licht and Sczepanski (2024) apply a similar span-labeling approach to British parliamentary debates, quantifying how often politicians mention particular social groups. These studies demonstrate the feasibility of automatic group mention detection, but remain monolingual, single-label, and limited to explicit group mentions.
@@
Beyond mentions, understanding how groups are portrayed is crucial. The Us vs. Them corpus (Huguet Cabot et al., 2021) annotates Reddit comments for target group, stance (supportive vs. hostile), and emotion, using a multi-task RoBERTa model. Card et al. (2022) study 140 years of U.S. congressional speeches on immigration, combining sentiment classification with custom frame lexica to reveal evolving and polarized frames. The Media Frames Corpus (Card et al., 2015) provides frame annotations (including cultural identity frames) for news articles. These resources focus on single-label stance or framing, and predominantly English texts.
@@
Separate but related is the task of inferring latent political ideology. Thomas et al. (2006) use SVMs on floor debate transcripts to classify support/opposition. Iyyer et al. (2014) develop neural networks to predict left/right alignment of speeches. More recent prompt-based methods with LLMs (Mishra et al., 2023) achieve zero-shot ideological scoring. While these approaches infer broad ideological leanings, they do not detect explicit identity mentions or support multilabel identity categories.
In-group/out-group rhetoric.
Discourse-level signals such as pronoun clusivity and coded language reveal populist identity appeals. Rehbein and Ruppenhofer (2022) annotate inclusive vs. exclusive uses of “we” in German parliamentary debates and train classifiers to disambiguate referents, highlighting us-versus-them framing. Mendelsohn et al. (2023) curate a dogwhistle lexicon and show that pretrained models struggle to detect covert slurs without knowledge grounding. These works emphasize the need for high-precision, context-aware annotation, but do not cover multilabel identity categories across many classes.
@@
Current identity-focused corpora are typically single-label, English-centric, and restricted to broad political categories or explicitly stated group mentions. They rarely offer the fine-grained, multilabel annotations needed to capture the complexity of real-world political speech, and they do not cover non-Latin scripts or languages such as Hebrew. We introduce the first publicly available Hebrew dataset for social identity detection, comprising 5,536 sentences from politicians’ Facebook posts annotated with twelve distinct identities—each grounded in survey-measured salience and expert-defined categories. By providing multilabel annotations across a rich set of ideologically, religiously, and socio-economically relevant identities, our resource enables more nuanced analyses of how political actors deploy identity language than has previously been possible in any non-English setting.
@@ 3.1 Panel Survey
In order to choose social identities for annotation, we utilized a combination of experts and survey instruments. We utilized a representative 12-wave panel survey of the Jewish population in Israel (N = 1,769), conducted between January 2019 and April 2021 (ANONYMIZED REFERENCE). 222Similar to other panel surveys in which local sample vendors were unable to include re-interview samples of Palestinian citizens in sufficient numbers (e.g., (Gidron et al., 2022), the panel survey did not include this population, reflecting a known deficiency in Israel’s survey sampling market. This time period contains four Israeli elections held in quick succession, providing a unique opportunity to investigate political dynamics in a condensed time frame. In 12 survey waves, respondents were asked to select up to three identities they identify with, out of a list of 28. These 28 identities were chosen by a panel of experts in Israeli politics as reflecting a broad spectrum of prevalent social identities, spanning ideological, value-based, religious, national, socio-demographic, and economic dimensions (e.g., Conservative, Leftist, Nationalist, LGBTQ+. See Appendix A for a full list). For inclusion in the textual annotation, we selected the 12 identities that emerged as most salient in the panel survey, each consistently surpassing a 5% selection threshold in the first five waves. These identities are Rightist, Leftist, Conservative, Liberal, Socially-oriented, Capitalist, Zionist, Palestinian, Honest, Security-oriented, Ultra-orthodox, and Democrat.333The 5% criterion was jointly applied to identities that respondents identify with and identities they disapproved of in a separate survey item.
@@ 3.2 Annotation Scheme
Below is a summary of the identity definitions used in the annotation scheme, accompanied by abridged examples translated from Hebrew (full definitions appear in appendix B).
@@ 4 Dataset
To sample sentences for annotation, we compiled a corpus of Facebook posts by Israeli members of parliament, political parties, and viable party candidates, posted in the equivalent timeframe of the survey (Dec. 2018 to Apr. 2021). This corpus comprises 64,174 posts containing 375,718 sentences. Our annotated dataset contains 5,536 Hebrew-language sentences sampled from the Facebook posts corpus. Each sentence is annotated with 12 binary annotations according to whether it expresses a given identity. The dataset is split into a training set (70%), validation set (15%) and test set (15%).
@@ 4.1 Inter-coder Reliability
The sentences were annotated by two of the authors. We evaluate inter-coder reliability on a sample of 304 sentences. The results, given in Table 1, indicate a mean Cohen’s Kappa of 0.77. Disagreements were resolved via consultation with all authors.
| Category | Percent Agreement | Cohen’s Kappa |
@@ 5 Modeling and Experiments
We model the task of identifying social identities from text using three types of training setups:
Multilabel encoder-models are fine-tuned to jointly learn the 12 identity labels. Single-label encoder models are fine-tuned separately for each label. For this training, we utilize the best performing base model in the multilabel encoder setup. Finally, Seq2seq LLMs are fine-tuned to generate a comma-separated list of all applicable Hebrew language labels for the given sentence.
We examine a set of encoder models and LLM decoders in the 2B-9B parameter ranges. For encoders, we use the multilingual mBERT (Devlin et al., 2019), and Hebrew-targeted encoders AlephBERT (Seker et al., 2022), HeRO (Hebron et al., 2023) and the base and large variants of DictaBERT (Shmidman et al., 2023). All encoder models are trained for up to 10 epochs with three learning rate (1e-5, 3e-5, 5e-5), and three type of loss (default, positive weight and focal loss), choosing the best performing checkpoint on the validation set.
For decoder LLMs, we use the multilingual Gemma 2 in the 2B and 9B variants (Team et al., 2024), the multilingual Qwen3-8B (Team, 2025), and the Hebrew-targeted DictaLM2.0 (Shmidman et al., 2024). All decoders are fine-tuned with QLORA (Dettmers et al., 2023) for up to 5 epochs with two learning rates (1e-5, 1e-4).
@@
| Gemma-9B | 0.717 | 0.698 | 0.705 | 0.645 | 0.759 | 0.684 | 0.753 | 0.591 | 0.662 | 0.749 | 0.758 | 0.671 | 0.737 | 0.804 | 0.650 | ||||||||||||||||
| Gemma-2B | 0.620 | 0.631 | 0.624 | 0.560 | 0.723 | 0.575 | 0.757 | 0.385 | 0.634 | 0.673 | 0.716 | 0.609 | 0.615 | 0.680 | 0.557 | ||||||||||||||||
| Qwen-8B | 0.665 | 0.463 | 0.542 | 0.605 | 0.508 | 0.308 | 0.700 | 0.410 | 0.541 | 0.568 | 0.625 | 0.516 | 0.515 | 0.650 | 0.554 | ||||||||||||||||
@@
| Single-label encoders | |||||||||||||||||||||||||||||||
| Dictabert-Large | 0.643 | 0.689 | 0.659 | 0.672 | 0.659 | 0.529 | 0.731 | 0.508 | 0.708 | 0.694 | 0.783 | 0.615 | 0.636 | 0.721 | 0.652 | ||||||||||||||||
Table 4 reports test set results for trained models. We observe that the most performant are seq2seq models, achieving the highest per-label F1 scores on all but one label. DictaLM2.0, specifically, achieves the best macro P/R/F1 scores, and best per-label scores on 8/12 identities. The performance of this model, continuously pre-trained from Mistral 7B on 100B Hebrew tokens, underscores the dependence identity detection on language specific cultural and political context.
We also note that two of the multilabel encoders, and , perform better on average than the single-label encoders finetuned for each identity separately using as a base model. achieved an score of 0.678 in a multi-label setting, compared to a mean of 0.659 when combining the single-label models. This indicates the benefit of joint learning and the interrelated nature of identities in this task.
@@ 5.1 Generalization to Parliamentary Speech
We examine whether the resulting classifier generalizes well to the parliamentary speech of the Israeli parliament (Knesset) floor by utilizing the IsraParlTweet dataset (Mor-Lan et al., 2024). We subset speeches from the equivalent time period of the Facebook data, classify all sentences with the best performing DictaLM2.0 model, and sample 500 of the sentences for human annotation. We perform a precision-oriented test by oversampling positive predictions for each identity. The results show a macro score of 0.72, on par with the Facebook test set, indicating the generalization ability of the model to Knesset data (for full results, see Table 5).

### v2

@@ Abstract
Political language is deeply intertwined with social identities. While social identities are often shaped by specific cultural contexts, existing NLP datasets are predominantly English-centric and focus on coarse-grained identity categories. We introduce HebID, the first multilabel Hebrew corpus for social identity detection. The corpus contains 5,536 sentences from Israeli politicians’ Facebook posts (Dec 2018–Apr 2021), with each sentence manually annotated for twelve nuanced social identities (e.g., Rightist, Ultra-Orthodox, Socially-oriented) selected based on their salience in national survey data. We benchmark multilabel and single-label encoders alongside 2B–9B-parameter decoder LLMs, finding that Hebrew-tuned LLMs provide the best results (macro- = 0.74). We apply our classifier to politicians’ Facebook posts and parliamentary speeches, evaluating differences in popularity, temporal trends, clustering patterns, and gender-related variations in identity expression. We utilize identity choices from a national public survey, comparing the identities portrayed in elite discourse with those prioritized by the public. HebID provides a comprehensive foundation for studying social identities in Hebrew and can serve as a model for similar research in other non-English political contexts.111https://github.com/guymorlan/hebid/
@@ 1 Introduction
Social identities—such as political ideology, religious affiliation, or demographic group membership—are powerful drivers of political behavior and public discourse (Tajfel and Turner, 1979). Yet existing NLP resources for identity detection remain almost entirely English‐focused, single‐label, and rely on coarse-grained categories (e.g., party or ethnicity). In this work, we introduce HebID, the first publicly released Hebrew dataset for fine‐grained, multilabel social identity detection in political text, grounded in both domain expertise and large‐scale survey evidence.
We frame the task as a sentence-level classification problem: the goal is to determine which, if any, of twelve social identities are being positively expressed or invoked within a given sentence.
The choice of social identities is empirically grounded in survey data. We utilized 12 waves of a national survey (N = 1,769) each of which included questions designed to identify the social identities most salient to Israeli citizens (e.g., Rightist, Ultra-Orthodox, Socially-oriented). These survey‐derived categories were then used to guide the annotation of 5,536 sentences sampled from Facebook posts by Israeli politicians (December 2018–April 2021), ensuring that our labels reflect real‐world identity salience.
We benchmark a suite of modern architectures on the Facebook data—(i) multilabel encoder models; (ii) single‐label encoder models; and (iii) 2B–9B‐parameter decoder LLMs, finding that Hebrew‐tuned decoder models (specifically, DictaLM2.0) achieve the highest macro‐F1 (0.743), outperforming encoder‐only baselines by over six points. We then assess cross‐genre generalization by applying our best model to 500 parliamentary speech excerpts, achieving a comparable macro‐F1 of 0.72.
Finally, we link three complementary sources in our analysis of how identities behave: politicians’ Facebook posts, parliamentary speeches from the Israeli parliament (Knesset), and survey responses that reflect public identification. We (1) compare identity prevalence and correlations between social media, parliamentary speech, and the public; (2) document how identity discourse surges around election cycles; (3) uncover coherent “bundles” of co‐occurring identities; and (4) quantify gender‐related variation in identity expression.
@@
Dataset and survey-grounded methodology: We introduce a framework for creating identity corpora by linking survey data to text annotation, and release HebID, the resulting Hebrew dataset.
Comprehensive benchmarks: We evaluate a range of encoder-only and decoder LLMs, establishing strong baselines for the task.
Cross-genre evaluation: We demonstrate our best model’s generalization to parliamentary speech, confirming its robustness.
External validation: We validate our classifier against an external expert survey (CHES-Israel), showing its alignment with party policy positions.
Sociolinguistic analysis: We use the classifier to reveal novel insights into identity dynamics across social media, parliamentary speeches, and public surveys.
Taken together, these contributions offer value on three fronts. For computational social scientists, HebID provides a novel tool for analyzing political discourse. For the NLP community, our work establishes a challenging new benchmark for multilabel classification in a low-resource language. Finally, our survey-grounded annotation framework serves as a methodological template for developing similar culturally-aware resources in other non-English contexts.
@@ 2 Related Work
Early work extended named‐entity recognition to capture social groups as entities. Zanotto et al. (2024) introduce GRIT, annotating Italian news and parliamentary text for spans referring to demographic, national, or partisan groups, and fine‐tune BERT to identify them. Licht and Sczepanski (2024) apply a similar span‐labeling approach to British parliamentary debates, quantifying how often politicians mention particular social groups. These studies demonstrate the feasibility of automatic group mention detection, but remain single-label and limited to explicit group mentions.
@@
Beyond mentions, understanding how groups are portrayed is crucial. The Us vs. Them corpus (Huguet Cabot et al., 2021) annotates Reddit comments for target group, stance (supportive vs. hostile), and emotion, using a multi‐task RoBERTa model. Card et al. (2022) study 140 years of U.S. congressional speeches on immigration, combining sentiment classification with custom frame lexica to reveal evolving and polarized frames. The Media Frames Corpus (Card et al., 2015) provides frame annotations (including cultural identity frames) for news articles. These resources focus on single‐label stance or framing, and predominantly English texts.
@@
Separate but related is the task of inferring latent political ideology. Thomas et al. (2006) use SVMs on floor debate transcripts to classify support/opposition. Iyyer et al. (2014) develop neural networks to predict left/right alignment of speeches. More recent prompt‐based methods with LLMs (Mishra et al., 2023) achieve zero‐shot ideological scoring. While these approaches infer broad ideological leanings, they do not detect explicit identity mentions or support multilabel identity categories.
In‐group/out‐group rhetoric.
Discourse‐level signals such as pronoun clusivity and coded language reveal populist identity appeals. Rehbein and Ruppenhofer (2022) annotate inclusive vs. exclusive uses of “we” in German parliamentary debates and train classifiers to disambiguate referents, highlighting us‐versus‐them framing. Mendelsohn et al. (2023) curate a dogwhistle lexicon and show that pretrained models struggle to detect covert slurs without knowledge grounding. These works emphasize the need for high‐precision, context‐aware annotation, but do not cover multilabel identity categories across many classes.
@@
Current identity‐focused corpora are typically single‐label, English‐centric, and restricted to broad political categories or explicitly stated group mentions. They rarely offer the fine‐grained, multilabel annotations needed to capture the complexity of real‐world political speech, and they do not cover non‐Latin scripts or languages such as Hebrew. We introduce the first publicly available Hebrew dataset for social identity detection, comprising 5,536 sentences from politicians’ Facebook posts annotated with twelve distinct identities—each grounded in survey‐measured salience and expert‐defined categories. By providing multilabel annotations across a rich set of ideologically, religiously, and socio‐economically relevant identities, our resource enables more nuanced analyses of how political actors deploy identity language than has previously been possible in any non‐English setting.
@@ 3.1 Panel Survey
To select social identities for annotation, we utilized a combination of experts and survey instruments. We utilized a representative 12-wave panel survey of the Jewish population in Israel (N = 1,769), conducted between January 2019 and April 2021 (Dvir-Gvirsman et al. 2022; see also Rivlin-Angert et al. 2025).222Similar to other panel surveys in which local sample vendors were unable to include re-interview samples of Palestinian citizens in sufficient numbers (e.g., Gidron et al. (2022)), the panel survey did not include this population, reflecting a known deficiency in Israel’s survey sampling market. This time period contains four Israeli elections held in quick succession, providing a unique opportunity to investigate political dynamics in a condensed time frame. In 12 survey waves, respondents were asked to select up to three identities they identify with, out of a list of 28. These 28 identities were chosen by a panel of experts in Israeli politics as reflecting a broad spectrum of prevalent social identities, spanning ideological, value-based, religious, national, socio-demographic, and economic dimensions (e.g., Conservative, Leftist, Nationalist, LGBTQ+. See appendix A for a full list).
For inclusion in the textual annotation, we selected the 12 identities that emerged as most salient in the panel survey, each consistently surpassing a 5% selection threshold in the first five waves. These identities are Rightist, Leftist, Conservative, Liberal, Socially-oriented, Capitalist, Zionist, Palestinian, Honest, Security-oriented, Ultra-orthodox, and Democrat.333The 5% criterion was jointly applied to identities that respondents identify with and identities they disapproved of in a separate survey item. For more information on the survey methodology, see appendix B.
@@ 3.2 Annotation Scheme
Below is a summary of the identity definitions used in the annotation scheme, accompanied by examples from the dataset translated from Hebrew (full definitions appear in appendix C). Some examples were abridged for brevity.
@@
3.3 Co-occurring Identities
Our annotation scheme is multi-label, allowing a single sentence to express several social identities simultaneously. Each annotated sentence is evaluated with respect to all included identities. For example, the following sentence expresses Democrat, Leftist, and Honest social identities:
[…]
Similarly, this sentence expresses both a Rightist and a Liberal identity:
[…]
@@ 4 Dataset
To sample sentences for annotation, we compiled a corpus of Facebook posts by Israeli members of parliament, political parties, and viable party candidates, posted during the same timeframe as the survey (Dec. 2018 to Apr. 2021). This corpus comprises 64K posts containing 375K sentences. Our annotated dataset contains 5,536 Hebrew-language sentences sampled from 4.8K unique Facebook posts. Each sentence is annotated with 12 binary annotations according to whether it expresses a given identity. The dataset is split into a training set (70%), validation set (15%) and test set (15%).
@@ 4.1 Inter-coder Reliability
The sentences were annotated by two of the authors. We evaluate inter-coder reliability on a sample of 304 sentences. The results, given in Table 1, indicate a mean Cohen’s of 0.77. Disagreements were resolved via consultation with all authors.
| Category | Percent Agreement | Cohen’s |
@@ 5 Modeling and Experiments
We model the task of identifying social identities from text using four types of training setups:
First, we establish Baseline models by training separate Logistic Regression and LinearSVC classifiers for each label using TF-IDF features. Then, Multilabel encoder-models are fine-tuned to jointly learn the 12 identity labels. Single-label encoder models are fine-tuned separately for each label. For this training, we utilize the best performing base model in the multilabel encoder setup. Finally, Decoder LLMs are fine-tuned to generate a comma-separated list of all applicable Hebrew language labels for the given sentence.
We examine a set of encoder models and LLM decoders in the 2B–9B parameter range. For encoders, we use the multilingual mBERT (Devlin et al., 2019), and Hebrew-targeted encoders AlephBERT (Seker et al., 2022), HeRO (Hebron et al., 2023) and the base and large variants of DictaBERT (Shmidman et al., 2023). All encoder models are trained for up to 10 epochs with three learning rates (1e-5, 3e-5, 5e-5), and three types of loss (default, positive weight and focal loss), choosing the best performing checkpoint on the validation set.
For decoder LLMs, we use Gemma 2 in the 2B and 9B variants (Gemma Team, 2024), Qwen3-8B (Qwen Team, 2025), and the Hebrew-targeted DictaLM2.0 (Shmidman et al., 2024). All decoders are fine-tuned with QLORA (Dettmers et al., 2023) for up to 5 epochs with two learning rates (1e-5, 1e-4).
@@
| Gemma-2‐9B | 0.717 | 0.698 | 0.705 | 0.645 | 0.759 | 0.684 | 0.753 | 0.591 | 0.662 | 0.749 | 0.758 | 0.671 | 0.737 | 0.804 | 0.650 | ||||||||||||||||
| Gemma-2‐2B | 0.620 | 0.631 | 0.624 | 0.560 | 0.723 | 0.575 | 0.757 | 0.385 | 0.634 | 0.673 | 0.716 | 0.609 | 0.615 | 0.680 | 0.557 | ||||||||||||||||
| Qwen‐8B | 0.665 | 0.463 | 0.542 | 0.605 | 0.508 | 0.308 | 0.700 | 0.410 | 0.541 | 0.568 | 0.625 | 0.516 | 0.515 | 0.650 | 0.554 | ||||||||||||||||
@@
| Single‐label encoders | |||||||||||||||||||||||||||||||
| DictaBERTLarge | 0.643 | 0.689 | 0.659 | 0.672 | 0.659 | 0.529 | 0.731 | 0.508 | 0.708 | 0.694 | 0.783 | 0.615 | 0.636 | 0.721 | 0.652 | ||||||||||||||||
| Baselines | |||||||||||||||||||||||||||||||
| LinearSVC | 0.398 | 0.339 | 0.361 | 0.400 | 0.170 | 0.300 | 0.490 | 0.230 | 0.430 | 0.330 | 0.520 | 0.390 | 0.390 | 0.350 | 0.350 | ||||||||||||||||
| LogisticRegression | 0.469 | 0.268 | 0.334 | 0.350 | 0.000 | 0.280 | 0.500 | 0.280 | 0.410 | 0.340 | 0.470 | 0.380 | 0.350 | 0.360 | 0.290 | ||||||||||||||||
Table 4 reports test set results for trained models. We observe that the most performant are decoder models, achieving the highest per-label F1 scores on all but one label. DictaLM2.0, specifically, achieves the best macro P/R/F1 scores, and best per-label scores on 8/12 identities. The performance of this model, continuously pre-trained from Mistral 7B on 100B Hebrew tokens, underscores the dependence identity detection on language specific cultural and political context.
We also note that two of the multilabel encoders, DictaBERTLarge and DictaBERT, perform better on average than the single-label encoders fine-tuned for each identity separately using DictaBERTLarge as a base model. DictaBERTLarge achieved an score of 0.678 in a multi-label setting, compared to a mean of 0.659 when combining the single-label models. This indicates the benefit of joint learning and the interrelated nature of identities in this task.
@@ 5.1 Generalization to Parliamentary Speech
We examine whether the resulting classifier generalizes well to the parliamentary speech in the Knesset by utilizing the IsraParlTweet dataset (Mor-Lan et al., 2024). We subset speeches from the equivalent time period of the Facebook data, classify all sentences with the best performing DictaLM2.0 model, and sample 500 of the sentences for human annotation. We perform a precision-oriented test by oversampling positive predictions for each identity. The results show a macro score of 0.72, on par with the Facebook test set, indicating the model’s ability to generalize to Knesset data (for full results, see Table 5).
5.2 External Validity
We examine the external validity of the classifier by correlating social identity discourse with external party policy rankings from the CHES-Israel expert survey (Zur and Bakker, 2025), which assigns parties scores on several policy domains. For each of the 16 political parties covered by both the expert survey and our Facebook corpus in year 2021, we calculate identity discourse scores, defined as the share of sentences posted by the party’s politicians which are classified as expressing said identity.
We match closely related policy issues from the expert survey to each social identity. For example, policy issue Economic Stance is linked with both Capitalist and Socially-oriented social identities. For better comparability, in some cases we additionally construct social identity dimensions by subtracting party means for two polar identities (e.g. , ). We then compute correlations between the CHES party policy scores and the predicted social identity discourse scores, for the 16 political parties.
Our analysis reveals a strong alignment with the expert-coded CHES data. Out of 21 correlations, 16 are statistically significant at and 13 are significant at . Of those 13, all fall within a strong effect size range, with absolute correlation values from to . For full results, see Table 7 in the appendix.